jhonyy9 (jhonyy9):

@ganeshie8 - just bc. this is a very very interesting problem - suppose 0/0 is defined fraction just in this case so how is this possibile ? 0/0 =2

jhonyy9 (jhonyy9):

let 0/0 being just in this case defined fraction 0=100-100 100-100 10^2 -10^2 -------- = -------------- = 100-100 10*10 -10*10 (10-10)(10+10) (10+10) 20 = ---------------- = ---------- = ---- = 2 10(10-10) 10 10 so in this way 0/0 = 2

jhonyy9 (jhonyy9):

@agent0smith @.Sam. @Directrix @TheSmartOne @Kainui

OpenStudy (aaronandyson):

Can we just make assumptions like that and then prove it?

OpenStudy (mww):

This is the same as the proof of 1 = 2 Let x = y = 1 Then x^2 = x = 1 x^2 - 1 = x - 1 (x+1)(x-1) = x - 1 x+ 1 = (x-1)/(x-1) = 1 Therefore 2 = 1 Of course this never works since you cannot divide by zero.

OpenStudy (welshfella):

you can make a number equal to any number if you allow 0/0 5 * 0 = 0 ---> 0/0 = 5 8 *0 = 0 ----> 0/0 = 8 so 5 = 8

OpenStudy (inkyvoyd):

there's a reason 0/0 is an indeterminate form ;)